OpenStudy (anonymous):

A particle is moving with the given data. Find the position of the particle. a(t) = t - 2 s(0) = 3 v(0) = 2

OpenStudy (anonymous):

integral of a(t) is t^2/2-2t+C, well, C=v(0)=2 integral of v(t) is t^3/6-t^2+2t+C, and C=s(0)=3 so s(t)=t^3/6-t^2+2t+3

OpenStudy (anonymous):

Thanks :)